Let a,b,c,d be positive rationals such that a+√b=c+√d then prove that a=c and b=d or b and d are square of rationals

Grade:10

1 Answers

Arun
6 years ago
Dear Sunny
 
Given a + √b = c + √d
Case (i): Let a=c
⇒ a + √b = c + √d becomes
a + √b = a + √d
⇒ √b = √d
∴ b = d
Case (ii): Let a ≠ c
Let us take a = c + k where k is a rational number not equal to zero.
⇒ a + √b = c + √d becomes
(c + k) + √b = c + √d
⇒ k + √b = √d
Let us now square on both the sides,
⇒ (k + √b)2 = (√d)2
⇒ k2 + b + 2k√b = d
⇒ 2k√b = d – k2 – b
 
sqrt(b) = [d - k² - b]/2k
 
Notice that RHS is a rational number.
 
Hence sqrt(b) is a rational number.
This is possible when b is a square of rational number.
 
Thus d is also a square of rational number  as
 
k + sqrt(b) = sqrt (d)
 
 
Regards
Arun (askIITians forum expert)